OpenStudy (anonymous):

if a and b are integers and 14/3 equals a/b what is the value of a plus b

OpenStudy (tkhunny):

Proportion? \(\dfrac{a}{b} = \dfrac{14}{3}\) Then: \(3a = 14b\) I suspect there are many more solutions than just one.

OpenStudy (anonymous):

The choices are 68, 66, 63, 60, 58

OpenStudy (anonymous):

Wats a good way to solve anyway

OpenStudy (snowfire):

You take the value of 14+3=17, and any multiples of that will work with the ratio

OpenStudy (tkhunny):

We want a+b. We have 3a = 14b Then 3a + b = 15b Then a + b = 15b - 2a That doesn't seem to be going anywhere. Let's just ry some.58-68 isn't very far. We have 3a = 14b a = 14 and b = 3 is a relatively obvious solution. No? 14 + 3 = 17 a = 28 and b = 6 is a relatively obvious solution. Just double both a and b. 28 + 6 = 34 a = 56 and b = 12 is a relatively obvious solution. Just double both a and b. 56 + 12 = 68 -- I think we have a winner!

OpenStudy (anonymous):

Thank u

OpenStudy (anonymous):

Perfect

OpenStudy (anonymous):

Why 4?

OpenStudy (tkhunny):

4 - because that's what worked.